Grade 12th passGeneral Physics

A block of mass 10 kg is moving in x direction with constant speed of 10 m/s. It is subjected to retarding force f= -0.1x joules / metre during its travel from x=20metres to x=30metres it's final kinetic energy will be

shubham
8 Years agoGrade 12th pass

1 Answer

DIWAKAR PANDEY
ApprovedApproved Tutor Answer8 Years ago
By energy conservation Ek1=1/2×10×100=500 After motion PE=fdx (-ve sign is only for retardation PE=0.1x2/2PE =25 By Energy conservation500=25+Ek(final)Ek(final)=500-25 Ek(final)=475 J